OpenStudy (anonymous):

Find the Exact Value of each expression (answer in Radians) A) sin^-1(1/2) B)cos^-1(1/2)

OpenStudy (e.mccormick):

Know your unit circle?

OpenStudy (anonymous):

I never learned it.

OpenStudy (e.mccormick):

OK. Well, you need to work on that. There is a way to remember it pretty easy using just the numbers 1, 2, and 3. For now, just google one up to use. If you think about it, \(\sin^{-1}\) or arcsin and \(\cos^{-1}\) or arccos have a very clear, reverse relationship to the non-arc versions.

OpenStudy (e.mccormick):

Lets say I have some angle, \(\theta\) and from it I get a value through the use of cos. That is an x value. So \(\cos\theta = x\) How do we get from x to \(\theta\)? With arccos. \(\cos\theta = x\) Take the arccos of both sides: \(\arccos(\cos\theta) = \arccos (x)\) The arccos of cos cancels and: \(\theta = \arccos (x)\) So, if you have a unit circle, you can look for x=1/2 and what \(\theta\) that works with. Then look for y=1/2 for the arcsin one.